a woman asked 10 boys to clean a 5 roomed house, they finished it in 30 minutes. How long will it take 15 boys to finish cleaning the house????

Answer by Theo(13342)

rate per person times number of people times time = quantity
r*p*t = q

quantity is 1 clean house.
t = 30 minutes
p = 10
you are solving for rate per person.

formula becomes:

r*10*30 = 1

divide both sides of the equation by 10*30 to get:

r = 1 / (10 * 30) = 1/300.

1 person can clean 1/300 of the house in 1 minute.

1/300 * 10 * 30 = 300 / 300 = 1 cleaned house, so the value of r is confirmed to be good.

r is the amount of the house tht 1 person can clean in one minute.

you have established that one person can clean 1/300 of the house in 1 minute.

now you want to know how long it would take 15 persons to clean the house.

the formula of r*p*t = q is used again, and becomes:

(1/300) * 15 * t = 1

now you want to solve for t.

you get t = (1 / 15) * 300 = 20 minutes.

10 boys can clean the house in 30 minutes.
15 boys can clean the house in 20 minutes, working at the same rate for each boy.

1/300 * 10 * 30 = 1

1/300 * 15 * 20 = 1

solution is it will take 15 boys 20 minutes to clean the house.